OpenStudy (anonymous):

Select the points that lie on the function h(x) = 3x 2. (1, 3) (1, 9) (-1, -3) (-1, 3)

OpenStudy (helder_edwin):

\[ \large h(x)=3x^2 \] right??

OpenStudy (anonymous):

Yes

OpenStudy (helder_edwin):

the first one for instance (1,3) u got x=1 and y=3 so \[ \large y=h(1)=3\cdot1^2=3 \] so this point DOES lie on the function h

OpenStudy (anonymous):

Oh ok..so it's the first one then?

OpenStudy (helder_edwin):

check the others in this way

OpenStudy (anonymous):

ok i will

OpenStudy (helder_edwin):

what i did proves the 2nd is not. the square cancels signs, so this rules out the 3rd. what about the 4th???

OpenStudy (anonymous):

the forth is one, it's just neg.

OpenStudy (helder_edwin):

\[ \large y=h(-1)=3\cdot(-1)^2= \]